Which sneakers are prefered?
female- running 12 cross training 28
Male- running 42 cross training 14
3. P(female | running)
4. P(female | cross training)
5. P(male | running)
6. P(male | cross training)
For questions 7 and 8, find the mean, median, and mode of each data set.
8.) 5 7 10 10 7 9 5 6 7 7 4


10.) 12 13 14 9 7 8 15 10 11 6

Answer by stanbon(75887)

Which sneakers are prefered?
female- running 12 cross training 28
Male- running 42 cross training 14
---------------------------------------------------
3. P(female | running) = 12/54
4. P(female | cross training) = 28/42
5. P(male | running) = 42/54
6. P(male | cross training) = 14/42
==========================================
Cheers,
Stan H.